# Multi-Agent Reinforcement Learning with Control-Theoretic Safety Guarantees for Dynamic Network Bridging

### v1

@@ IV Safe Tracking Control
Agent Type Safety Mechanism Explicit Penalty Truncation Avg Coverage Avg Safety Violations
Baseline A No No No 41% 13,880
Baseline B No Yes No 0.01% 0
Approach A Yes No No 39% 0
Approach B Yes Yes No 16% 0
Approach C Yes Yes Yes 22% 0
TABLE I: Comparison of agent performance, safety mechanisms, explicit penalties, and safety violations.
V Neural Network Architecture
We employ the neural network architecture for action-value estimation presented in [7], which is tailored to address the challenges of dynamic network bridging in a cooperative multi-agent setting, enabling the agents to extrapolate information about spatial and temporal dependencies while collaborating effectively. Such architecture combines Graph Neural Networks (GNNs) for relational modeling and exploits their message-passing mechanisms compatibly with the task optimized in the environment [Ref Section]. Furthermore, a Long Short-Term Memory (LSTM) is employed to cope with temporal dependencies and the partial observability of the task. The model is trained in a Centralized Training Decentralized Execution (CTDE) [15] fashion, where the agents optimize the same action-value function parameterization during training. However, during execution, each agent acts based on its local observations and the shared representations from its neighbors, adhering to the decentralized nature of the task.
We continue by describing the key components of the architecture below.
V-1 Graph Attention Layers
Graph Convolutional Reinforcement Learning (DGN) [16] facilitates cooperation between the agents by enabling them to share their latent representations with their immediate neighbors. To this end, we leverage Graph Attention Network (GAT) [17] to capture the spatial and relational dependencies between agents in the network. Specifically, we employ two Multi-Headed GAT that enable each agent to attend to and integrate information from its neighboring agents adaptively. This encoding allows the agents to condition their actions based on the dynamic network topology and inter-agent relationships effectively.
Additionally, we integrate target entities into the sharing process by equipping them with this same encoding module to produce their latent representation. If an agent’s neighborhood includes target entities, the agent can gather their representations and condition its actions accordingly.
V-2 Long Short-Term Memory
To handle the partial observability and temporal dynamics of the environment, we integrate a LSTM [18] layer into our architecture. The LSTM layer maintains a memory of past observations, allowing the agents to make informed decisions based on past interactions, partial observability, and the evolving environment.
During training, we employ observation stacking, aggregating observations over multiple time steps. Such a process gives the LSTM layer the necessary temporal context for effective learning in the presence of partial observability.
V-3 Dueling Action Decoder
To estimate action values effectively, we incorporate a Dueling Action Decoder [19] in our architecture. This component comprises separate streams for estimating state values and advantages, which are then combined to produce the final action value estimates.
VI Experimental Setup
We evaluated our approach in a simulated 2D environment with three agents and two moving targets. The environment was normalized with axes ranging from 0 to 1. Agents and targets had a communication range of 0.25, and their movement offset was set to 0.05 for the calculation of the next target point $w$. The physics simulator allowed us to simulate the non-linear Unmanned Aerial Vehicle (UAV) dynamics while we updated the positions of the agents in their paths toward the decided target point.
Our multi-agent strategies were trained across 1M training steps. Training episodes are truncated and Partial Episode Boostraping [20] was performed after 100 steps taken by each agent. Agents were initialized at fixed starting positions that satisfy safety condition (11), and the moving targets were randomly placed at a certain distance apart to ensure realistic difficulty and a feasible connection. Target motion and placing follow a seeded random pattern to control the scenarios during training and testing. The primary focus of our experiments was to demonstrate the effectiveness of our safe tracking control system in enabling the MARL approach to learn policies that satisfy safety constraints. To establish the importance of this contribution, we designed a set of experiments with varying training configurations:
  • Baseline A: We first trained a typical MARL approach without any safety tracking control, serving as a baseline.
  • Baseline B: To investigate the impact of explicit penalization for safety violations, we trained the same MARL approach as in Baseline A but with a high penalty (-100) imposed whenever an agent violated the safety constraints.
  • Approach A: This configuration represents our key contribution, where we trained the agents with the safe tracking control system enabled. Agents employed Algorithm 1 to update their target points while respecting the safety conditions.
The results from Approach A demonstrated the effectiveness of our safe tracking control system in enabling the agents to learn policies that accomplish the task while adhering to safety constraints.
To further solidify our approach and highlight its advantages, we conducted additional experiments:
  • Approach B: Here, we combined the safe tracking control system with explicit penalization for safety violations, similar to Baseline B. This configuration aimed to analyze the agents’ behavior when both safety measures and penalization for safety violations were employed simultaneously.
  • Approach C: Building upon Approach B, we additionally introduced episode termination whenever safety constraints were violated. This experiment assessed the impact of a more stringent safety enforcement mechanism.
If enabled, the safe tracking control system was used to prevent agents from continuing their movement if they violated the safety condition (11), such as going too close to other agents and risking collision. During training, in case the safety condition was violated, the agents involved were forced to stop their movement and would remain blocked until a new action toward safe directions was computed by at least one of the agents involved. To evaluate our agents, we ran 100 evaluation episodes disabling the safety tracking control system and counting the number of times that safety conditions were violated for every position update (14).
VII Results
Table I shows the results of our experiments evaluating different learned MARL strategies. We report the average communication coverage achieved by the agents as well as the number of safety violations observed when the safe tracking control system was disabled during evaluation.
Baseline A agents, trained without any safety mechanisms, achieved 41% coverage but incurred 13,880 safety violations on average. Introducing an explicit penalty of -100 for violating safety constraints (Baseline B) reduced the coverage to only 0.01% but did successfully prevent any safety violations.
In our Approach A, agents achieved 39% coverage (2% less than Baseline A) while completely avoiding any safety violations. This demonstrates that safe tracking control allows the agents to learn policies that respect safety constraints without significantly compromising task performance and without the need for explicit constraint violation penalties. Adding an explicit penalty for safety violations to the safe tracking control (Approach B) reduced coverage to 16% while still avoiding violations. Further adding episode truncation when violating safety (Approach C) increased coverage to 22% while maintaining zero violations.
VIII Discussion
Our work presents a novel hybrid approach that combines MARL with control-theoretic methods to address a complex cooperative task while ensuring safety guarantees. The theoretical analysis provides three key results. First, we prove that each agent’s setpoint update only affects the safety of its one-hop neighboring agents, enabling decentralized coordination. Second, we establish that Algorithm 1 guarantees the preservation of the safety condition for all agents under specific assumptions. Third, we derive an analytical condition to efficiently evaluate potential safety violations during setpoint updates.
The experimental results highlight the importance of our hybrid approach. Agents trained without any safety mechanisms (Baseline A) achieved high task coverage but incurred numerous safety violations. Introducing an explicit penalty for violating safety constraints (Baseline B) successfully prevented violations but at the cost of significantly compromised task performance. In contrast, our Approach A, which incorporates the safe tracking control system based on Algorithm 1, allowed agents to learn policies that respect safety constraints while maintaining reasonable task coverage, without explicit constraint violation penalties.
Despite these strengths, there are some limitations to our current work. While we have demonstrated the approach’s effectiveness on the network bridging task, we have not extensively evaluated its scalability to significantly larger swarm sizes or other domains. Moving forward we plan to investigate the scalability and adaptability of our approach across different domains and with varying swarm sizes.

### v2

@@ III Safety Filter for Tracking Control
Training Conditions Evaluation Results
Safety Filter Explicit Penalty Truncation Avg Coverage Avg Safety Interventions
Baseline No No No $(37.63\pm 16.62)\%$ $26.99\pm 12.31$
Approach A Yes No No $(50.14\pm 12.72)\%$ $20.45\pm 7.49$
Approach B Yes Yes No $(45.02\pm 11.81)\%$ $15.25\pm 7.82$
Approach C Yes Yes Yes $(10.86\pm 6.44)\%$ $9.76\pm 5.72$
TABLE I: Comparison of the trained agents performance across 100 evaluation episodes.
IV Decentralized Safety-Aware Agents
This section details the components of our approach to ensure safe interactions while learning effective positioning strategies in networked and decentralized multi-agent tasks, by combining MARL with our safety filter mechanism. The core intuition is to employ Algorithm 1 since the training phase and inform the agents of eventual safety filter activations.
IV-A Guiding learning via decentralized safety filters
By enforcing adherence to explicit safety conditions defined in (7), agents experience realistic operational constraints and encounter possible deadlock scenarios whenever direct trajectories toward their target points trigger the safety filter. These constraints inherently shape the policy optimization landscape during training: 1) creating deadlock conditions agents might be blocked by their respective safety filters unless they reconsider their actions; 2) avoiding unsafe states of the environment, which are not going to be visited by the agents. Consequently, the agents need to leverage interactions and communication with neighboring agents to collaboratively avoid or resolve deadlock situations by generating alternative, safety-compliant paths to achieve their objectives effectively.
IV-B Safety-informed edges
Leveraging the graph nature of the task introduced in Section II-C, we explicitly introduce edge-level features linked to the conditions determining the activation of the decentralized safety filter. Specifically, edge features comprise base attributes controlling the activation of the safety filter and a binary safety activation indicator between connected agents. Formally, given two connected agents $i$ and $j$, the edge features are defined as:
$f_{l}^{ij}=[\mathbf{f},S_{i,j}],$ (16)
where $\mathbf{f}$ denotes a vector of available attributes related to the activation of the decentralized safety filter, and the binary indicator $S_{i,j}$ explicitly signals whether the safety filter has been activated between the two agents, thereby preventing setpoint update.
IV-C Policy parameterization
To demonstrate our approach we train the agents with a simple independent Q-Learning algorithm [16] and employ the same GNN architecture proposed in [8] for action-value estimation. The architecture integrates Graph Attention Network (GAT) [17] and Long Short-Term Memory (LSTM) layers [18], enabling effective modeling of spatial and temporal dependencies essential for collaborative decision-making. The GAT layers allow agents to dynamically attend to and integrate neighbor information, significantly improving inter-agent coordination by adaptively sharing latent representations through message passing. The LSTM captures temporal dependencies and partial observability, allowing agents to leverage historical context in their decisions. The architecture is trained using a centralized training decentralized execution (CTDE) framework, where the agents optimize the same action-value function parameterization during training while acting independently based on local observations and the interaction with the other agents.
V Experimental Setup
In this section, we describe in detail the specific environment implementation and training configurations employed for our MARL agents.
V-A Dynamic network bridging environment implementation
Environment: The simulated environment is normalized to a $[0,1]\times[0,1]$ grid, though agents can move beyond these limits. Each entity has a communication radius of 0.20, which defines the area within which it can detect and communicate with others.
Agent observations: At each time step, agents receive an observation consisting of their local one-hop graph structure. This includes node-level features for themselves and all directly connected neighbors, as well as edge-level features for each neighbor pair. An example is shown in Figure 3, where Agent A observes Agent B. The observation includes both their node features $f_{n}^{A},f_{n}^{B}$ and the corresponding edge features $f_{i}^{AB}$ from a successful connection. Node-level features include the agent’s ID, current coordinates, last action taken, and the positions of the targets A and B. These features provide basic local context for decision-making. Edge-level features reflect spatial and safety-related information, as introduced in Section IV-B. For each neighboring agent $j$, agent $i$ observes relative position offsets $\Delta p_{x}=p_{x,i}-p_{x,j}$, $\Delta p_{y}=p_{y,i}-p_{y,j}$, and their Euclidean distance $d_{i,j}=\sqrt{\Delta p_{x}^{2}+\Delta p_{y}^{2}}$. Including safety-informed edge features allows agents to adjust their actions based not only on position but also on the state of the safety filter. This supports more informed and coordinated decision-making during training and execution.
Agent actions: At the beginning of every time-step, agents make decentralized decisions, generating target points $w$, utilizing their local observation to choose the direction of their next movement. This is encoded in a two-dimensional action space, with each dimension having three options: move forward, move backward, or hold along the corresponding $p_{x}$ or $p_{y}$ axis. Given the direction chosen by the agents, a new target point $x_{sp}$ is calculated for each agent using a fixed offset equal for both the $p_{x}$ and $p_{y}$ axes.
Reward: Agents are rewarded using a reward function that motivates agents to form a connected network bridging the targets. Such reward signal combines three components:
$R_{\text{base}}(s)=\frac{|C_{\text{max}}(s)|}{|\mathcal{V}|},$ (17)
The overall reward combines these three elements:
Simulated UAV dynamics: For our experiments, we use a six-degrees-of-freedom (6-DOF) unmanned aerial vehicle (UAV) as an agent, where its dynamics, control laws, and safety sets used in our simulations are defined in the Appendix of [7]. For the UAV simulations, the nonlinear model of the drone is used.
Target mobility: To increase task complexity, each target moves deterministically toward waypoints located outside the communication range of any agent. When a target reaches a waypoint, it pauses for a fixed number of time-steps (5 in our implementation) before selecting a new one. Waypoint selection ensures that at least two agents (and up to $N$) are required to maintain a communication bridge between the targets.
Figure 3: A typical configuration of our system. Agents are represented with letters A, B, C and targets with T. Agent A and B have established a successful connection. Agent A receives node features $f_{n}^{A}$, $f_{n}^{A}$ and edge features $f_{n}^{AB}$ as inputs.
V-B Training configurations
We train our agents using episodes of 100 time steps, for a total of one million steps. For generalization, we initialize agent positions, target locations, and target motion patterns using different seeded random states across training and testing scenarios. During evaluation, each trained agent is tested in 100 new episodes, and we measure performance metrics with the safety filter consistently enabled.
We systematically evaluated different training configurations to assess the impact of our safety mechanism:
  • Baseline: A vanilla MARL approach with no safety filter enabled. Edge features include only the base features $\mathbf{f}$.
  • Approach A (Safety Filter): Our main contribution. Agents use Algorithm 1 to prevent safety infringements. The binary safety activation indicator is incorporated into the agents’ edge-level features ($\mathbf{f}$ and $S_{i,j}$).
  • Approach B (Safety + Penalty): We add an explicit penalty of $-10$ to Approach A whenever an agent attempts an action that would trigger the safety filter.
  • Approach C (Safety + Penalty + Early Truncation): Building on Approach B, we truncate episodes immediately upon safety-triggering situations, thus enforcing even stricter safety adherence.
Every approach adopts the same learning algorithm and GNN architecture. In any configuration employing the safety filter, an agent violating (7) has its trajectory toward $x_{sp}$ blocked at its current position.
Furthermore, we train an alternative of Approach A, called Approach A-Node, to investigate the benefits of our proposed integration of safety-informed edges IV-B. In this alternative approach, the agents utilize only the node-level features with the addition of a binary indicator $\mathbb{S}$ in case the agent had his safety filter activated during the previous step.
VI Results
(a) Mean return achieved during training. Approach A (with safety-informed edges) converges to higher rewards compared to Approach A-Node.
(b) Average number of safety filter activations per episode. Approach A consistently triggers fewer interventions.
Figure 4: Impact of safety-informed edge features.
Table I summarizes our experimental evaluation of various learned MARL strategies under different training conditions. This comparison was conducted to evaluate the impact of our safety filter. We report the average communication coverage achieved by the agents and the average number of times the safety filter was activated during evaluation.
Furthermore, Figure 4 focuses on a comparison between Approach A and Approach A-Node in terms of the mean return, i.e. the sum of the rewards, achieved by the agents during training and the mean number of safety activations triggered by the agents.
VI-A Impact of the safety filter
Baseline agents, trained without any safety mechanisms, achieved an average coverage of $(37.63\pm 16.62)\%$ but required frequent interventions from the safety filter, averaging $26.99\pm 12.31$ triggers per episode. In contrast, our Approach A, which incorporates our safety filter during training without explicit penalties, significantly improved average coverage to $(50.14\pm 12.72)\%$ and reduced safety filter activations to $20.45\pm 7.49$.
Introducing explicit penalties alongside the safety mechanism (Approach B) maintained competitive coverage at $(45.02\pm 11.81)\%$ and further reduced the frequency of safety filter interventions to $15.25\pm 7.82$. However, implementing truncation upon safety violations (Approach C) considerably decreased task performance to $(10.86\pm 6.44)\%$, though it also minimized the safety triggers to $9.76\pm 5.72$.
These results demonstrate that integrating our safety filter during training and including safety-informed edges allows agents to learn better strategies for the dynamic network bridging environment.
Results on Approach B and Approach C highlight how the introduction of explicit penalties might reduce safety interventions, which could have practical benefits in specific applications. However, this led to a decrease in task performance, which was conspicuous for Approach C, suggesting that agents might exploit the activation of the safety filter to achieve better performance.
VI-B Impact of the safety-informed edges
Figure 5: Sequence of actions executed by agents trained with Approach A when deployed in a LVC environment.
Figure 4(a) shows the mean return (cumulative reward) curves for Approach A and Approach A-Node, over the course of training. Approach A (with safety edge features) converges to higher returns compared to Approach A-Node, indicating that safety-informed edges facilitate more effective coordination strategies. Additionally, Figure 4(b) presents the average frequency of safety filter activations across training episodes. The results show that Approach A triggers fewer safety interventions than Approach A-Node, particularly in the later stages of training.
These findings suggest that edge-level safety features, rather than node-level, help achieve better performance while incurring in less safety activations. Figure 5 shows an example of a sequence of actions operated by our decentralized agents trained with Approach A when deployed in a live, virtual, and constructive (LVC) environment.
VII Discussion
Our findings demonstrate that integrating a decentralized safety filter with MARL provides a versatile solution for multi-agent networked tasks, where agents can directly communicate with their one-hop neighbors. By embedding the safety mechanism in each agent, we ensure local checks on potential collisions, thereby containing safety concerns within these immediate communication ranges. This design preserves the decentralized nature of our approach while providing global safety assurances. We applied the framework to the dynamic network bridging task, which we extended to include more realistic target mobility and non-linear UAV dynamics.
To support coordination, we included safety-informed edge features in the agents’ observations. These features indicate when the safety filter is active, allowing agents to adjust their actions based on local constraints. This helped agents avoid collisions, maintain communication links, and resolve potential deadlock situations through coordination.
Empirical evaluations confirm that our approach effectively improves task performance while assuring safety. Although adding explicit penalties or early-termination strategies reduced safety activations, these methods often compromised the agent’s utility. In contrast, our decentralized filter and safety-informed edges guide safe exploration without requiring extensive reward engineering or outcome truncations. It is important to highlight that the design principles generalize naturally to larger numbers of agents and targets and it can be applied to different environments than the one under study here. For example, we can apply similar safety filters to other networked navigation tasks or expand the existing safety filter to include complex metrics.
Looking ahead, we plan to investigate scalability for swarms of varying sizes, refine message-passing protocols for more sophisticated risk management, and evaluate domains with adversarial behaviors or dynamic constraints. Overall, our results highlight the potential of combining local safety controllers with MARL in networked multi-agent settings, underscoring that edge-level safety signals are crucial for guiding agents toward globally safe and high-performing outcomes.